OpenStudy (anonymous):

I REALLY REALLY NEED HELP....:) can anyone help me solve this problem is a linear equation with parentheses...the problem is: 3x-4(2x+2)=2......o yeah show your work please......THANKS

hero (hero):

3x-8x-8 = 2 -5x-8 = 2 -5x = 10 x = -2

OpenStudy (anonymous):

-2